Block A , which has a weight of 10N, stays at rest the initial position the spring which has a stiffness of K=200N/m, is compressed by 10 cm and the released . if the coefficient of kinetic friction between the block and the surface is Block A , which has a weight of 10N, stays at rest the initial position the spring which has a stiffness of K=200N/m, is compressed by 10 cm and the released . if the coefficient of kinetic friction between the block and the surface is nBlock A , which has a weight of 10N, stays at rest the initial position the spring which has a stiffness of K=200N/m, is compressed by 10 cm and the released . if the coefficient of kinetic friction between the block and the surface is µk=0.2, determine the acceleration of the block just after the spring is released
1
Expert's answer
2020-04-27T10:14:40-0400

As per the question,

Weight of the block "(w)=10N"

"m=10\/9.80=1.02kg"

Stiffness of the spring "(K)=200N\/m"

Compression in the spring "(x)=10cm= 0.1m"

Coefficient of kinetic friction"(\mu_k)=0.2"

let the acceleration of the block after the release is a,

So, "Kx-\mu w=ma"

"\Rightarrow 200\times 0.1-0.2\times 10=1.02\times a"

"\Rightarrow a=\dfrac{18}{1.02}=17.64 m/sec^2"